OpenStudy (anonymous):

I'm trying to figure this out, but I can't seem to get it right. It says:

OpenStudy (anonymous):

Which of the following is the solution of log x+3 (.001)=-3

OpenStudy (anonymous):

I tried putting it in exponential form, and looked it up, but every time I change it to exponential form, I get: (x+3)^-3=.001, and I'm pretty sure that's not right. How do I do this? It's getting confusing.

OpenStudy (unklerhaukus):

is it \[\log(x+3)^{.001}=-3\]?

OpenStudy (unklerhaukus):

Use \[\large\log_bx=y\quad\iff\quad x=b^y\]

OpenStudy (anonymous):

Ummm... the x+3 looks like a base, and .001 isn't an exponent, so I don't think so

OpenStudy (unklerhaukus):

oh , like this? \[\log_{x+3}(.001)=−3\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Also, I've seen a similar formula you gave, only it looks like this: y=b^x >> x=log b (y). What's the difference between the two?

OpenStudy (unklerhaukus):

um there is no difference i just choose x where they chose y and visa versa. \[\large y=b^x\quad \iff\quad \log_b (y)=x\] hmm, can you express .001 as a power of ten?

OpenStudy (anonymous):

That would be 10^-2, right?

OpenStudy (unklerhaukus):

.1 is 10^-1 .01 is 10^-2 so .001 is 10 ^-...

OpenStudy (anonymous):

Oh 10^-3. Whoops. Okay, got it.

OpenStudy (unklerhaukus):

Right , so \[\log_{x+3}(.001)=−3\\\log_{x+3}(10^{-3})=−3\]

OpenStudy (anonymous):

So... 10^-3=(x+3)^-3 ?

OpenStudy (unklerhaukus):

yes!

OpenStudy (anonymous):

So when I'm solving this, can I cancel the exponents out, or would that be bad?

OpenStudy (unklerhaukus):

that's correct the exponents on both sides are the same so they can cancel 10 = x+3

OpenStudy (anonymous):

So x=7. Okay, got it. Thank you for helping me!! That really fixed my confusion.

OpenStudy (unklerhaukus):

yay you got it !